OpenStudy (anonymous):

7^5 ÷ 7 ^2 = 343 a) True b) False

OpenStudy (anonymous):

a. true

OpenStudy (anonymous):

7^5=16807 7^2=49 16807/49=343 a)true

OpenStudy (radar):

\[7^{5}\div7^{2}=7^{(5-2)}=7^{3}=343\]